OpenStudy (anonymous):

Hello can someone show me/explain how to do this problem step by step please! Ty

OpenStudy (usman1995):

where's the problem?

OpenStudy (anonymous):

#18

OpenStudy (unklerhaukus):

18.\[\begin{align} 4-(2x-1)&=3(2x-1)\\ %4+(-1)\times(2x)+(-1)\times(-1)&=3\times(2x)+3\times(-1)\\ 4-2x+1&=6x-3\\ &= \end{align}\] .... keep going

OpenStudy (usman1995):

these really are simplest questions, see we can't show u each and every step as it's against openstudy policy, because it donot encourage students, here's what u can do solve the questions as the unkel has shown u and send in those pics i'll see for each step if they are right.. okay?

OpenStudy (anonymous):

4-(2x-1)= 3(2x-1) 4-2x+1=6x-3 4+3+1=6x+2x 8 = 8x x=1